# captionUnsupported document class

### v1

@@ V-D Safe Planning in Real Crowd Environment
Full trajectory knowledge. First, we assume that the robot knows the trajectories of all humans in the field for the entire horizon. In this setting, the coefficients of CBF, CLF, and Distance reward were set to $0.3$, $0.01$, and $0.3$, respectively to emphasize safety constraints since the number of obstacles increases and the complexity of movement becomes more challenging for ensuring safety. Since the simulations are discrete, the CBF reward cannot completely guarantee safety. As a result, it is necessary to introduce a buffer between the barrier radius and collision radius, set to $1\ m$ and $0.7\ m$, respectively. This buffer accounts for the inherent limitations of discrete simulations and provides a margin of safety. We included CoBL-Diffusion- (CoBL-) without covariance weight in our evaluation to investigate the effect of the weight on covariance (19) in encouraging convergence of the earlier control sequence. The results of each model are summarized in Table I. For reference, the actual 200 pedestrian data with $3-8$ $m$ movements are shown as Human. The high collision rate of Human is attributed to the defined collision radius being larger than the actual one.
CoBL demonstrated its ability to generate safe plans even in complex environments with multiple moving people. On the other hand, CBF-QP is unable to avoid moving obstacles in a dynamic multi-agent environment. Although the trajectories generated by VO can safely navigate around multiple obstacles, they often lack smoothness and may become impractical or unpredictable for surrounding humans. Similar to simulations in the single-agent environment, using dCoBL or DCoL rewards makes it challenging to ensure safety. Comparing CoBL with CoB, while equivalent quality trajectories were generated in the simple environment, the contribution of CLF reward to goal-reaching is more prominent in this more complex setting; the trajectory is constantly modified by the CBF reward, leading to deviations from the goal without the CLF reward. When comparing the CoBL with the CoBL-, it seems that the temporal weighting introduced for the covariance (19) to promote convergence of the earlier control sequence slightly reduces the collision rate; these results prompt further investigation to better understand the impact of the temporal weighting.
@@
FullTraj $1$ $0.41\pm 0.20$ $0.07\pm 0.03$
FourSecTraj $0.5$ $0.40\pm 0.20$ $0.07\pm 0.03$
TwoSecTraj $4$ $0.36\pm 0.20$ $0.07\pm 0.03$
InitTraj $19.5$ $0.45\pm 0.30$ $0.07\pm 0.03$

### v2

@@ V-D Safe Planning in Real Crowd Environment
Full trajectory knowledge. First, we assume that the robot knows the trajectories of all humans in the field for the entire horizon. In this setting, the coefficients of CBF, CLF, and Distance reward were set to $0.3$, $0.01$, and $0.3$, respectively to emphasize safety constraints since the number of obstacles increases and the complexity of movement becomes more challenging for ensuring safety. The CBF barrier radius was set to $1\ m$, providing a buffer for a collision radius of $0.7\ m$. We included CoBL-Diffusion- (CoBL-) without covariance weight in our evaluation to investigate the effect of the weight on covariance (19) in encouraging convergence of the earlier control sequence. The results of each model are summarized in Table I. For reference, the actual 200 pedestrian data with $3-8$ $m$ movements are shown as Human. The high collision rate of Human is attributed to the defined collision radius being larger than the actual one.
CoBL demonstrated its ability to generate safe plans even in complex environments with multiple moving people, with an average generation time of $4.8$ seconds. On the other hand, CBF-QP is unable to avoid moving obstacles in a dynamic multi-agent environment. Although the trajectories generated by VO can safely navigate around multiple obstacles, they often lack smoothness and may become impractical or unpredictable for surrounding humans. The result of the Diffuser has the worst distance from the goal because they are modified by the CBF-QP after generating a plan. In contrast, our model integrates collision avoidance into the planning process, enabling the synthesis of safe trajectories from the outset. Similar to simulations in the single-agent environment, using dCoBL or DCoL rewards makes it challenging to ensure safety. Comparing CoBL with CoB, while equivalent quality trajectories were generated in the simple environment, the contribution of CLF reward to goal-reaching is more prominent in this more complex setting; the trajectory is constantly modified by the CBF reward, leading to deviations from the goal without the CLF reward. When comparing the CoBL with the CoBL-, it seems that the temporal weighting introduced for the covariance (19) to promote convergence of the earlier control sequence slightly reduces the collision rate; these results prompt further investigation to better understand the impact of the temporal weighting.
@@
FullTraj $1$ $0.41\pm 0.20$ $0.74\pm 0.34$
FourSecTraj $0.5$ $0.40\pm 0.20$ $0.68\pm 0.29$
TwoSecTraj $4$ $0.36\pm 0.20$ $0.67\pm 0.32$
InitTraj $19.5$ $0.45\pm 0.30$ $0.67\pm 0.32$